### v2

# ROS2WASM: Bringing the Robot Operating System to the Web

Abstract
The Robot Operating System (ROS) has become the de facto standard middleware in robotics, widely adopted across domains ranging from education to industrial applications. The RoboStack distribution, a conda-based packaging system for ROS, has extended ROS’s accessibility by facilitating installation across all major operating systems and architectures, integrating seamlessly with scientific tools such as PyTorch and Open3D. This paper presents ROS2WASM, a novel integration of RoboStack with WebAssembly, enabling the execution of ROS 2 and its associated software directly within web browsers, without requiring local installations. ROS2WASM significantly enhances the reproducibility and shareability of research, lowers barriers to robotics education, and leverages WebAssembly’s robust security framework to protect against malicious code. We detail our methodology for cross-compiling ROS 2 packages into WebAssembly, the development of a specialized middleware for ROS 2 communication within browsers, and the implementation of www.ros2wasm.dev, a web platform enabling users to interact with ROS 2 environments. Additionally, we extend support to the Robotics Toolbox for Python and adapt its Swift simulator for browser compatibility. Our work paves the way for unprecedented accessibility in robotics, offering scalable, secure, and reproducible environments that have the potential to transform educational and research paradigms.
I Introduction
The Robot Operating System (ROS) has emerged as the de facto standard robotics middleware, facilitating communication between multiple nodes in a network [1]. Its extensive libraries and active community support have led to widespread adoption across domains ranging from education to industrial applications [2]. However, as the ROS user base expands, issues related to the accessibility of ROS distributions have surfaced. ROS2WASM addresses these concerns by bridging the gap between high-performance scientific computing and web-based applications, making advanced robotics tools more accessible to a broader audience. To aid readers in navigating the technical terms and concepts used throughout this paper, a glossary is provided in Table I.
ROS 2 @ localhost
Figure 1: Execution of a ROS 2 publisher and subscriber setup within a Jupyter Notebook environment running entirely in the browser, with no local installation required (www.ros2wasm.dev). The notebook creates publishers for topics topic_a and topic_b, as well as two minimal subscribers listening to these topics. The log output at the bottom shows the concurrent execution of these nodes, with messages being published and received, verifying seamless communication between ROS 2 nodes directly within the browser-based interface.
ROS 2 has made significant strides by enabling native installations on major platforms, including Ubuntu, Windows, and macOS. The RoboStack effort has further extended ROS 2’s reach by allowing installations within conda environments [3]. Nevertheless, variations in package versions across platforms can complicate the replication of workspaces, posing challenges for both educational and research applications [4].
Replicating workspaces is particularly beneficial in scenarios such as teaching robotics courses [5] and reproducing research papers [6]. Instructors often struggle to provide students with the necessary tools, resorting to maintaining a computer lab or providing instructions for workplace replication. However, students’ diverse computing environments can lead to inconsistencies in how ROS packages function due to differences in operating systems, library versions, and system configurations. ROS2WASM mitigates these issues by providing a standardized execution environment within the browser itself. By compiling ROS 2 to WebAssembly, students can interact with identical software configurations regardless of their hardware or operating system, running ROS 2 directly in their browsers (e.g., Chrome, Firefox, Safari, Edge). ROS2WASM thus eliminates any local setup or reliance on university servers (Figure 1).
In research, reproducibility faces similar obstacles. Traditional environments are often rigid and difficult to adapt, and version compatibility issues can impede reproducibility efforts. While Docker containers provide a potential solution, installing Docker and managing containers can be daunting and time-consuming for many aspiring roboticists [7, 8].
We propose a solution based on WebAssembly (WASM) [9, 10], a low-level bytecode format designed to run efficiently in web browsers. WASM is fast, secure, portable, and well-sandboxed across all major browsers, making it ideal for high-performance web applications.
Instead of relying on Docker images, we package ROS 2 workspaces for web browsers, enabling easy distribution via simple links. These browser-based environments can run on any operating system and processor architecture, from smartphones to high-end desktops. By cross-compiling ROS packages to WebAssembly, ROS2WASM runs ROS nodes directly in the browser and facilitates communication between them. Unlike other ROS tools that depend on an existing ROS system or server connection, ROS2WASM creates a fully isolated, sandboxed environment within the browser.
Although ROS2WASM involves initial complexity in cross-compilation and middleware development, this complexity is largely encapsulated on the developer side. For end users, ROS2WASM offers a lower barrier to entry compared to Docker and similar solutions. ROS2WASM builds upon RoboStack [3], which integrates ROS with the Conda package manager. RoboStack offers several key advantages, including reproducible environments, robust version management, compatibility across platforms and Linux distributions, distribution of pre-built binaries, environment isolation, and simplified package creation.
Our contributions are as follows:
  1. 1.
    A custom ROS 2 middleware implementation, rmw-wasm, cross-compiled into WebAssembly modules, enabling communication between ROS 2 modules within browser environments. This middleware also enables communication with external physical robots running standard ROS 2 installations using https://wiki.ros.org/roslibjs.
  2. 2.
    Cross-compilation of numerous ROS 2 packages to WebAssembly, including rclpy, to provide Python support for beginners.
  3. 3.
    Deployment of www.ros2wasm.dev, which showcases an example browser environment for ROS2WASM.
  4. 4.
    Cross-compilation of the Robotics Toolbox for Python and adaptation of its Swift simulator for compatibility with WebAssembly, providing a concrete example of adaptation that can serve as a foundation for further extensions by other users.
TABLE I: Glossary
Term Description
Boa Build tool for conda packages.
Conda An open-source package management system and environment management system that allows users to install and manage packages, dependencies, and environments across different platforms.
Conda-forge A community-driven platform providing a vast collection of conda packages, built and maintained using continuous integration (CI) services to ensure cross-platform compatibility.
Cross-compiler A compiler that generates executable code for a platform different from the one on which the compiler is running.
DDS Data Distribution Service, a middleware protocol and API standard for data exchange between nodes in a distributed system, ensuring reliable, real-time communication.
Emscripten Compiler toolchain to convert C/C++ code into WebAssembly.
Emscripten-forge A repository of conda packages built using Emscripten, enabling the cross-compilation of C/C++ code to WebAssembly.
emscripten::val A utility provided by the Emscripten toolchain that facilitates the interaction between C++ and JavaScript by allowing C++ code to call and manipulate JavaScript objects and functions seamlessly within a WebAssembly environment.
JupyterLite A JupyterLab distribution which runs on a web browser.
RoboStack A project that integrates the Robot Operating System (ROS) with the conda package manager, enabling easy installation and management of ROS distributions across different platforms.
Toolchain A set of programming tools used together in a development process, typically including a compiler, linker, and debugger, among others, to automate the building and deployment of software.
WebAssembly (WASM) Low-level binary format designed for efficient execution on modern web browsers.
Web Worker A JavaScript API that allows running scripts in background threads, enabling concurrent execution without affecting the performance of the main browser thread.
II Related Works
This section overviews existing technologies related to WebAssembly, ROS in the browser, and browser-based robotics environments that contextualize our work.
II-A Overview of ROS 2 and RoboStack
Since its inception, ROS has been pivotal in advancing robotics research and development [11]. It has enabled the creation of complex robotic systems that can operate in diverse environments, from autonomous vehicles [12] to industrial automation [13]. The modular architecture of ROS, combined with its extensive library of packages, has made it a powerful tool for both prototyping and deploying robotic systems.
ROS 2 [1], the successor of ROS, builds on the strengths of its predecessor while addressing several limitations. It introduces a new communication middleware based on the Data Distribution Service (DDS), which provides real-time communication capabilities, improved security, and better support for multi-robot systems. DDS enables ROS 2 nodes to communicate over a network in a decentralized and scalable manner, making it ideal for distributed robotic applications.
RoboStack [3] is a project that integrates ROS with the Conda package manager, significantly enhancing the accessibility and usability of ROS across different platforms. By leveraging Conda, RoboStack allows users to install ROS distributions in isolated environments, ensuring reproducibility and compatibility across different operating systems and architectures. This approach simplifies the installation process, reduces dependency conflicts, and allows for easier management of ROS packages. RoboStack also facilitates the distribution of pre-built binaries, which can be cross-compiled for various platforms. This capability is central to the development of our proposed ROS2WASM, where we cross-compile ROS to WebAssembly to enable seamless deployment of ROS packages in a web environment.
II-B WebAssembly
WebAssembly (WASM) is a relatively new technology in the web development landscape [9], which emerged in 2017 based on a collaboration between major browser vendors – Chrome, Edge, Firefox, and Safari.
WASM is a binary code format designed for a stack-based virtual machine, optimized for performance, security, efficiency, and portability. It can be executed within web browsers as a complement to JavaScript or independently in other environments. These characteristics make WASM ideal for high-performance applications, including games, scientific visualizations, simulations, and developer tools. Due to its capabilities, WASM is particularly well-suited for implementing a robotics environment in the browser.
Early demonstrations of WASM’s potential included porting the game AngryBots, developed by Unity Technologies, to the web [14]. Similarly, Doom 3 was ported to WASM using Emscripten [15], showcasing the ability to run large, demanding C++ programs within a browser. WASM’s adaptability and performance make it appealing for robotics applications.
For ROS2WASM, we use Emscripten [16], an open-source compiler toolchain for converting C/C++ code to WebAssembly. Emscripten supports the compilation of various languages into WASM by leveraging LLVM [17], enabling their execution on the Web, Node.js, or other WASM runtimes. The pipeline for compiling C/C++ code to WASM with Emscripten is illustrated in Figure 2. Additionally, Emscripten can compile C/C++ runtimes of other languages such as Python, facilitating the indirect execution of those languages in the browser. Specifically, we build on top of emscripten-forge, a project that implements a build pipeline to compile Conda packages to Web Assembly. Emscripten-forge makes use of the same build tools as RoboStack, making it well suited to cross-compile ROS 2 to Web Assembly.
Figure 2: Overview of the pipeline for compiling C/C++ source code into WebAssembly (WASM) using Emscripten. The process begins with the C/C++ source code (e.g., talker.cpp), which is passed through the Emscripten compiler. Emscripten converts this source code into three primary outputs: a WebAssembly module (talker.wasm), an HTML document (talker.html) that serves as the interface to load and run the WASM module, and JavaScript glue code (talker.js) that manages the interaction between the WebAssembly module and the web environment. This pipeline allows native C/C++ applications to be executed in the browser, offering near-native performance and seamless integration with web technologies.
II-C ROS in the Browser
The closest analogue to ROS2WASM is the ROS on Web project [18], which shares similar goals of enabling ROS nodes to run entirely within a browser by cross-compiling C++ code to WebAssembly and using web workers for internal communication. However, a significant limitation of ROS on Web is its closed-source nature, which restricts access to the implementation details. While the project indicates that it replaces the ROS 2 middleware with a custom design and employs web workers, the exact methods remain undisclosed.
Another relevant project is the roswasm suite [19], an open-source collection of libraries designed to assist in cross-compiling C++ ROS nodes to WebAssembly. However, the roswasm suite is primarily focused on the outdated ROS 1 distribution, limiting its applicability to ROS 2 users.
II-D Browser-Based Robotics Environments
Several other projects, while not directly aligned with ROS2WASM’s objectives, are relevant to the broader context of developing robotics environments for web browsers. The Robot Web Tools website (https://robotwebtools.github.io) offers a comprehensive overview of open-source libraries and tools that facilitate the integration of robotic frameworks with the web. ROSWeb [20] aggregates all the ROS widgets provided by Robot Web Tools into a single web application, offering an interactive and user-friendly interface that requires minimal programming expertise.
One of the leading examples in this domain is Foxglove Studio [21], which focuses on the visualization and debugging of robotic tasks. Although Foxglove Studio’s primary focus is observability rather than direct robotic control, its cross-platform support for ROS 1 and ROS 2 and its availability as both a desktop application and a browser-based tool make it a significant contribution to the field. However, Foxglove Studio functions as an add-on to existing ROS installations rather than a standalone solution. Furthermore, Foxglove Studio is no longer open source and free for general use.
The rosbridge suite [22] enables non-ROS systems to communicate with ROS via JSON, supporting both ROS 1 and ROS 2 but requiring existing ROS installations, thus not being a complete alternative for running ROS in the browser. The ROS Control Center [23] is a web application that uses rosbridge to establish a WebSocket connection to a robotics system running ROS. In contrast, ROSboard [24] does not depend on the rosbridge suite. Instead, it implements a custom Tornado web server that functions as both a web server and a WebSocket server, providing a lightweight and streamlined interface for interacting with ROS systems.
III Methods and Demonstrations
This section details the methodology employed to develop ROS2WASM. We begin with the design of a custom middleware architecture for ROS 2 with WebAssembly, enabling communication within browser environments (Section III-A). Next, we describe the cross-compilation process extending RoboStack to support WebAssembly (Section III-B), followed by the web integration strategy using JupyterLite for browser-based access (Section III-C). Additionally, we cover adaptations to the Robotics Toolbox for Python and Swift Robotics Simulator for the sandboxed web environment (Section III-D). Finally, we discuss methods for communicating with non-ROS physical robots, demonstrating a practical application of ROS2WASM (Section III-E).
III-A Custom Middleware Architecture
Our custom middleware architecture comprises three main packages: rmw-wasm-cpp, wasm-cpp, and wasm-js (Figure 3). This design draws inspiration from rmw-email [25], which contains a middleware for sending and receiving strings over email, and an RMW implementation that allows ROS 2 to use this middleware for message exchange.
Figure 3: Architecture of the rmw-wasm middleware enabling ROS 2 communication in the browser. The system comprises three main components: rmw-wasm-cpp (interfaces ROS 2 packages with the WebAssembly-based middleware), wasm-cpp (the C++ core that processes messages and communicates with JavaScript), and wasm-js (manages network participants and message traffic within the browser). Communication between wasm-cpp and wasm-js is facilitated by emscripten::val, allowing seamless integration between C++ and JavaScript.
The rmw-wasm-cpp package acts as an adapter between ROS 2 packages and the middleware implementation; it contains all the functions required by the rmw interface. A primary function of this package is to convert binary ROS messages into YAML strings using dynamic message introspection (dynmsg). It then passes these strings to the wasm-cpp package for further processing. Conversely, when messages are retrieved by subscribers, rmw-wasm-cpp converts the YAML string messages back to binary ROS messages to return to the requester.
The middleware implementation is divided into wasm-cpp (implemented in C++) and wasm-js (implemented in JavaScript). Communication between these two packages is facilitated by emscripten::val, which allows JavaScript functions to be called from C++ and vice versa [16].
The wasm-cpp package serves as the core of the implementation and functions as a bridge to JavaScript modules (Figure 4). In this package, ROS elements are constructed hierarchically, with the smallest unit being a participant. Any ROS subscriber, publisher, service server, service client, action server, or action client is modeled as a participant with a specific role. Upon initialization, each participant receives a unique identifier for tracking purposes.
Publishers and subscribers are the simplest participant types and are responsible for sending and receiving messages, respectively. Service participants are derived from these basic types; service clients and servers consist of a publisher-subscriber pair (Figure 5). Before a service client can invoke a service, it verifies the availability of a service server through a query to wasm-js, which tracks all available entities. The client publishes a request to the service request topic and awaits the server’s response via the service response topic. The server publishes and subscribes to these topics in reverse. Figure 4 illustrates the relationships between the participants as implemented in the wasm-cpp package.
Participant name
role
gid
registration()
deregistration()
Publisher publish() Subscriber get_message() ServiceServer take_request()
send_response()
ServiceClient send_request()
take_response()
is_service_available()
Figure 4: Class diagram of the wasm-cpp package, showing the relationships between ROS 2 entities. The Participant class is the base, with Publisher, Subscriber, ServiceServer, and ServiceClient as derived classes. Each class handles specific ROS 2 functions, such as publishing, subscribing, and managing service requests and responses, enabling communication within the WebAssembly environment.
Service ClientPublisherSubscriberService ServerSubscriberPublisher/request     topic/response     topic
Figure 5: Data flow between a ROS service client and server within wasm-cpp. The service client initiates a request by publishing to the /request_topic, which is received by the service server’s subscriber. The server processes the request and then publishes a response to the /response_topic, which is received by the client’s subscriber. This setup demonstrates how service clients and servers in ROS are built on top of publisher-subscriber pairs, ensuring communication via specific request and response topics.
Messages are transmitted in YAML format from participants to wasm-js, which resides on the browser’s main thread (Figure 6). wasm-js is responsible for tracking network participants and managing message traffic. To support multiple ROS nodes in the browser, wasm-js spawns each node as a separate web worker. Web workers enable parallel execution of scripts in the background, allowing concurrent operations without affecting the performance of the main thread [26].
Once the web workers are operational, wasm-js stores all incoming data from publishers in message stacks. These message stacks act as temporary storage, holding messages until they are requested by subscribers. The stored data is then sent back to subscribers upon request, ensuring efficient communication between nodes.
ROS 2 rmw-wasm-cpp wasm-cpp wasm-js emscripten::val ROS message YAML string JSON object dynmsgjs-yaml
Figure 6: Message conversion flow within the rmw-wasm middleware. A ROS message is initially handled by rmw-wasm-cpp, where it is converted to a YAML string using dynamic message introspection (dynmsg). The YAML string is then passed to wasm-cpp, which bridges the communication to wasm-js via emscripten::val. In wasm-js, the YAML string is further converted into a JSON object using the js-yaml library. This flow ensures that ROS messages can be efficiently processed and transmitted within the browser-based WebAssembly environment.
III-B Cross-compilation
The cross-compilation process leverages the RoboStack ecosystem [3], extending its tools for WebAssembly. The vinca tool, which retrieves package information such as versions and dependencies from rosdistro and generates build recipes for the boa build tool, was adapted to support the WASM target platform. This modification was relatively straightforward due to vinca’s existing support for multiple platforms, requiring only minimal changes to the build script generation.
All ROS 2 packages must be recompiled specifically for WebAssembly to function in the ROS2WASM environment. For external libraries and dependencies, our approach builds upon the https://emscripten-forge.org, which is planned to be incorporated into the conda-forge ecosystem in the future. Emscripten-forge already provides many commonly used external libraries (including Python packages and numerical toolboxes) pre-compiled for WebAssembly.
As part of the development efforts for ROS2WASM, several dependencies were added to emscripten-forge, which sometimes necessitated the addition of patches tailored to WASM, e.g. introducing preprocessor guards for Emscripten. These patches are intended to guide future users on best practices and common pitfalls when targeting WebAssembly.
The overall build process closely follows that of RoboStack, with continuous integration set up to build and upload packages to an anaconda.org channel, available at https://github.com/RoboStack/ros-humble.
III-C Web Integration
After building individual packages, the next challenge is serving them on a website. Our approach enables two distinct types of web deployments: For applications compiled from C/C++ source code (as shown in Figure 2), we directly cross-compile the code to a WASM module and serve it as a standalone web application (Figure 8). For writing and executing ROS 2 applications directly in the browser, we utilize JupyterLite. By specifying the required packages in an environment.yaml file and running jupyter lite build, a complete website is generated, built on top of JupyterLite, with all specified packages installed. Users can then run commands such as import rclpy in a Python environment within the browser, using rclpy as they would in a typical ROS 2 setup (Figure 1). For those preferring C++, the rclcpp package can be used with the xeus-cpp kernel for JupyterLite.
III-D Robotics Toolbox for Python and Swift
This section briefly reviews how non-ROS robotics applications can be adapted for WASM. While the Robotics Toolbox for Python required no changes, adapting the graphics-based Swift Robotics Simulator [27] to the sandboxed web environment meant addressing limitations in threading, networking, and file system access. We cross-compiled Swift to WebAssembly using Emscripten and the PyJS framework, which required three major modifications.
Swift
Figure 7: Swift is a lightweight browser-based simulator built on top of the Robotics Toolbox for Python. This simulator provides robotics-specific functionality for rapid prototyping of algorithms, research, and education. Here, we introduce Swift running in the browser without any installation required: www.ros2wasm.dev.
First, we needed a new messaging system to replace WebSockets, which are not fully supported in WASM. This was achieved by passing messages via Python/Javascript interoperability using PyJS. Second, we had to find an alternative way to launch the Swift environment without starting an HTTP server. We eliminated the HTTP server by loading the web interface in an embedded iframe. Third, we needed to add hooks to load Robotics Toolbox assets from the packaged Python environment. This was achieved by intercepting asset fetching so that assets could be retrieved from the PyJS file system.
The resulting implementation demonstrates an approach for adapting existing projects to work within Emscripten and WASM constraints, which may inform future WebAssembly ports. Users can now interact with the simulator through a Python terminal embedded in a web page (Figure 7).
III-E Communication with a non-ROS Robot
To showcase the wide range of possible applications of ROS2WASM, we demonstrate the manipulation of a physical robot. Data can be sent and received from a robot with the publishers and subscribers developed in Section III-A, as long as there is a communication bridge between the robot and the web browser. There are two primary ways to communicate with a robot: Firstly, configure a robot to launch a rosbridge server to send information to the browser; however, this method requires the robot to be running ROS for creating a rosbridge. An alternative way for robots which do not run ROS natively is to create a bridge between the browser and the robot’s system. The bridge can be established via Bluetooth, WiFi, or through a wired connection. Here, we demonstrate ROS2WASM with the LEGO BOOST Vernie robot that is connected to any computer via Bluetooth. A pre-existing library, lego-boost-browser [28], is used to establish the interface between ROS and Vernie via a Web Bluetooth API.
LEGO Boost Vernie
Figure 8: Interface for controlling the LEGO BOOST Vernie robot using a ROS publisher to send LED color commands. The user can connect or disconnect from the robot via Bluetooth and use the interface to publish different LED colors by starting or stopping the “rainbow” publisher. The console logs below display the sequence of colors being published in real time, illustrating the robot’s response to the published messages.
The lego-boost-browser library can send commands to Vernie to change the LED color or activate the two motors for locomotion. We created a publisher that simply publishes different LED colors as strings from Vernie’s acceptable list of colors. When these messages reach the message stacks of the respective topic, they are forwarded to the robot through lego-boost-browser. Figure 8 shows how a web page running an LED color publisher appears, displaying the user interface with connect/disconnect buttons, publisher controls, and a console that logs the sequence of colors being published in real-time to demonstrate the robot’s response to commands. To control the position of the Vernie robot, we created a publisher that sends geometry_msgs/Twist messages that contain linear and angular velocity vectors to express the robot’s velocity in free space.
IV Conclusion
We demonstrated the feasibility and potential of running ROS 2 within a web browser using WebAssembly. Our work encompasses the design and implementation of three major components: a custom middleware architecture, a method for cross-compiling ROS 2 packages to WebAssembly, and the deployment of a web-based platform that allows users to interact with ROS environments directly in their browsers without the need for local installation or configuration.
Our website, www.ros2wasm.dev, enables the public to engage with a range of demonstrations that showcase the capabilities of ROS2WASM. These demonstrations vary in user interaction level and complexity, highlighting the versatility and scalability of the platform. By successfully running and facilitating communication between multiple ROS nodes within the browser, we have validated the core concept of ROS2WASM. This increased accessibility opens up new possibilities for education, research, and collaboration in robotics, making advanced tools more widely available. The current platform allows easy sharing of these environments via a simple link. However, further efforts are needed to simplify the setup process for novice users, particularly in creating and deploying their own ROS-based web applications.
ROS2WASM also has significant potential in industrial applications. Web-based ROS could enable robust server-based distributed control systems for multi-agent coordination, allowing operators to monitor and control robot fleets from any device without specialized software installations. Additionally, ROS2WASM could facilitate scalable and flexible cloud-based robotic processing, where computationally intensive tasks can be executed on remote servers while maintaining responsive browser-based interfaces on local devices.
Looking ahead, future work could explore integrating Zenoh as the middleware for ROS 2 [29]. Zenoh, a lightweight communication protocol recently added as an official RMW implementation for ROS 2, offers the potential to natively support WebAssembly. This could eliminate the need for our custom middleware, thereby streamlining the architecture and further improving performance and compatibility.
Another promising direction is adapting packages that contain graphical user interfaces such as the popular rviz, Gazebo, and MoveIt for browser compatibility. Such adaptations would further solidify the viability of running ROS 2 in the browser, which could lead to even broader adoption and innovation in web-based robotics.